imagen a table;
Fruit: Orange Apple Banana Strawberry other
Number of
students: 3 8 11 6 4

1) Suppose a student in the classroom is picked at random. Explain how to find the probability that the student's favorite fruit is strawberry.Express the probability in multiple ways.

2) Suppose a student in the classroom is picked at random. Explain how to find the probability that the student's favorite fruit is not an apple or a banana. Express the probability in multile ways.
-please explain throughly

Answer by edjones(8007)

3 + 8 + 11 + 6 + 4 =32
.
1) 6/32=3/16=.1875=18.75%
.
2) 8+11=19
32-19=13
13/32=.40625=40.625%
.
Ed